OpenStudy (anonymous):

Adult tickets to the fall play cost $8 and student tickets cost $4. The drama class sold 30 more adult tickets than student tickets to the fall play. If the class collected $840 from ticket sales, how many adult tickets were sold? The drama class sold _____ adult tickets.

OpenStudy (anonymous):

@olipop2002

OpenStudy (anonymous):

@mathmale

OpenStudy (zpupster):

Adult tickets =x student ticket = y so dollars *tickets = total sales first equation 8x+4y = 840 what else do we know? we know that 30 more adult tickets were sold adult tickets=student tickets + 30 or equation 2 x=y+30 or putting the variables on one side x-y=30 do you still need help?? can u get it from here?

OpenStudy (zpupster):

hello??

OpenStudy (anonymous):

yes

OpenStudy (zpupster):

can u get it from here? yes ok

OpenStudy (anonymous):

so?

OpenStudy (zpupster):

can u solve it now???

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

i feel stupid

OpenStudy (zpupster):

no do not feel stupid just something you have not learned yet

OpenStudy (anonymous):

it's like i have learned it, but over the break i forgot how to solve it.

OpenStudy (zpupster):

lets pick up where we left off we need to find one of the variables (x or y) and then we can get the other lets add the 2 equations together to eliminate a variable: 8x+4y = 840 x- y= 30 to eliminate a variable, lets multiply the 2nd eqaution by 4, you will see why. 8x+4y = 840 4x - 4y=120 now add and we remove y ________________ 12x + 0 = 960 getting x by itself we divide both sides by 12 12x/12 = 960/12 x=???

OpenStudy (anonymous):

80?

OpenStudy (zpupster):

correct --so we sold 80 adult tickets now we find y or student tickets by plugging x in the 2nd equation. 80 - y =30 add y both sides and also subtract 30 80 - 30 + y-y = 30 - 30 + y we get 50 = y 50 student tickets are we done no!! lets do a check by putting in our found values. 8(80) + 4(50) = 840 it checks go over the steps again and remember that you have 2 unknowns adult tickets and student tickets 2 unknowns 2 equations and so on.